Explain why there is no angle theta that satisfies cos =1/2 and sec=2 ?

Found 2 solutions by richard1234, Alan3354:
Answer by richard1234(7193)

You typed:

Explain why there is no angle theta that satisfies cos =1/2 and sec=2 ?

However cosine, secant, and all other trig functions require an argument such as theta or 2*theta. Therefore we cannot answer your question completely; please re-post. If you meant cos theta = 1/2 and sec theta = 2, there is a value of theta, e.g. 60 degrees or pi/3 radians.

Answer by Alan3354(69443)

Explain why there is no angle theta that satisfies cos =1/2 and sec=2 ?
========
There is.
cos(60) = 1/2
sec(60) = 2